# ShaRP: SHAllow-LayeR Pruning for Video Large Language Models Acceleration

### v1

@@ Abstract
Video Large Language Models (VLLMs) face the challenge of high computational load during the pre-filling stage due to the processing of an enormous number of visual tokens. Although attention-based pruning methods are widely used to accelerate inference, trials at early decoder layers often result in significant performance degradation, especially under high compression rates. We argue that while attention-based pruning inherently holds the potential to identify the most relevant visual tokens, its effectiveness in shallow decoder layers is limited by factors such as positional encoding bias and insufficient information interaction. In this paper, we propose an improved attention-based pruning framework, termed ShaRP, that integrates segment-aware causal masking, positional debiasing, and token deduplication for enhanced token selection. It enables effective pruning at shallow layers while maintaining stable performance under high compression rates without retraining. Extensive experiments demonstrate that ShaRP achieves competitive performance across multiple video understanding benchmarks, establishing a new paradigm for accelerating VLLM inference.
@@ 1 Introduction
Vision Large Language Models (VLLMs) [1, 2, 3, 4, 5] have brought deep video understanding task into a new era. They use a vision encoder to convert video streams into many tokens, which interact during pre-filling. Such a process thus incurs quadratic computational complexity relative to the token count. To mitigate this, numerous studies [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20] have focused on reducing visual token counts to compress information and accelerate VLLM inference.
Among various compression strategies, attention-based ones [6, 7, 8, 21, 14] offer a compelling approach. Unlike visual cue-guided pruning that operates purely on visual similarity [10, 22, 23, 24, 13], attention-based methods leverage the inherent cross-modal reasoning ability of the model to estimate token importance, inherently aligning the pruning process with semantic relevance. It not only enhances interpretability [25, 26, 27] but also enables adaptive compression that dynamically focuses on content most relevant to the textual query. Representative studies such as FastV [6] show that visual information becomes less critical in deeper decoder layers, motivating aggressive pruning early. However, this strategy struggles with high compression on long videos, as performance drops sharply, leading subsequent works [7, 8, 19] to prune more conservatively in deeper layers, losing much of the computational savings. Viewed differently, we advocate the potential of attention in shallow layers as an effective criterion for token selection. To explore this, we conduct a pilot study comparing attention scores across shallow and deep layers. As shown in Fig. 2, visual analysis reveals that visual tokens near the sequence end receive abnormally high attention in shallow layers. When properly tackled, shallow layer attention exhibits distributions similar to deep ones, enabling early yet reliable pruning.
@@
Our work is built upon a core insight: effective attention-based pruning in shallow layers is achievable when the three limitations above are properly addressed. To this end, we propose a novel training-free framework comprising three components — Segment-Aware Causal Masking (SegM), Positional Bias Calibration (PosC), and Register Token Deduplication (RegD) — for efficient shallow-layer token pruning in video understanding. First, to enhance information interaction, inspired by [34], which observed that locally focused heads in Transformers form near-diagonal block attention, segmenting complex reasoning into phrasal units. We adapt this to the visual domain by partitioning videos into content-aware segments, intensifying local interactions and alleviating attention collapse [30, 31]. Second, our method calibrates positional bias in attention scores through a rapid global bias estimation. Finally, to balance between the importance and redundancy, high-attention register tokens are refined through a three-step pipeline composed of pre-filtering, deduplication, and post-filling, which eliminates redundancies while enriching the representation with broader contextual information. Extensive experiments demonstrate that our approach achieves high compression in shallow LLM layers while maintaining state-of-the-art performance, establishing a new paradigm for attention-based token pruning.
@@ 2.1 Visual Cue-Guided Token Pruning
Since their emergence, VLLMs [35, 36, 37, 3, 38, 39, 4] have demonstrated significant potential across diverse applications, including VQA [40, 41, 42], video captioning [43, 44], contextual reasoning [45, 46, 47] etc.. Earlier approaches employ a limited number of learnable queries to aggregate visual information. For instance, VideoLLaMA [48] adopts a Q-Former [39] to distill dense visual features into compact query sets. While GlimpsePrune [49] introduces a trainable Visual Importance Predictor (VIP) within the VLLM to score and prune visual tokens at different layers. The recent method EPIC [50] adopts a Progressive Consistency Distillation strategy to re-finetune the model, enabling it to learn rich visual information even from incomplete visual inputs. Token pruning can be achieved in a training-free manner. Among these, LLaVA-PruMerge [9], VisionZip [10], and VisPruner [12] leverage the inherent sensitivity of vision encoders to visual tokens to identify and retain visually dominant tokens. DivPrune [22] formulates pruning as a Max-Min Diversity Problem (MMDP), ensuring that the set of retained visual tokens possesses maximum diversity. Nevertheless, the temporal redundancy inherent in video data calls for temporal-aware pruning mechanisms. DyCoke [13] merges tokens within fixed temporal windows. Other methods [23, 18, 19] remap the numerous visual tokens back to the frame level, enabling them to identify and merge highly similar visual content across frames while retaining unique content. Among these, LLaVA-Scissor [23] identifies non-overlapping visual regions via Semantic Connected Components (SCCs). FastVID [18] performs segmentation based on video content, combining temporal segmentation with spatio-temporal merging. HoliTom [19] utilizes dynamic programming to merge redundant information across adjacent frames.
@@ 4.2 Positional Bias Calibration
Positional Encodings (PE) equip VLLMs with token order information, whereas also introduce inherent positional bias [53] that favors tokens at specific locations. A typical example is the long-term decay issue [30] in RoPE (Eq. 2), where attention scores decrease as the relative token distance increases. While in the context of video token pruning, existing methods [6, 7, 8] usually identify more important visual tokens by referring attention scores from the last text token . They typically assume that visual tokens with higher scores are more relevant to the query. However, we argue that positional bias significantly distorts the distribution of attention scores, which may lead to misinterpretation of this mechanism. As shown in Fig. 4 (left-top), our analysis on the VideoMME [52] benchmark at 20% retention reveals that most top-ranked tokens (orange regions) cluster near the final frames, reflecting a clear positional bias rather than true semantic relevance. Fig. 4 (right-middle) further illustrates a case where the question targets a mid-video event, yet attention remains skewed toward later frames, leaving semantically important regions under-attended (blue regions).
@@ 4.3 Pruning and Register Token Deduplication
| Method |
MVBench ↑ |
MLVU ↑ | VideoMME ↑ | Average ↑ | |||||||||
| Overall | Short | Medium | Long | |||||||||||||
| Duration | 16s | 160min | 3120min | 160min | 13min | 330min | 3060min | Score | % | |||||||
| LLaVA-OneVision [3] | 100% | 40.8 | 57.9 | 56.4 | 63.2 | 58.6 | 70.3 | 56.6 | 48.8 | 59.0 | 100 | |||||
| FastV [6] | 25% | 11 | 55.3 | 55.1 | 60.3 | 56.5 | 68.2 | 55.0 | 46.3 | 56.8 | 96.2 | |||||
| PDrop [7] | 28.5% | 11 | 56.7 | 54.2 | 61.1 | 56.4 | 68.0 | 53.6 | 47.7 | 57.1 | 96.7 | |||||
| Feather [8] | 25% | 11 | 54.8 | 54.7 | 60.8 | 55.0 | 65.3 | 53.6 | 46.2 | 56.3 | 95.4 | |||||
| FrameFusion [24] | 31% | - | 56.7 | 55.8 | 62.2 | 57.4 | 68.1 | 55.4 | 48.6 | 57.8 | 97.9 | |||||
| ShaRP | 28.2% | 11 | 57.0 | 56.2 | 62.7 | 57.9 | 69.8 | 56.9 | 48.0 | 58.5 | 99.0 | |||||
| FastV [6] | 20% | 9.31 | 55.8 | 55.3 | 59.5 | 55.8 | 66.4 | 54.2 | 46.8 | 56.6 | 95.9 | |||||
| PDrop [7] | 19.6% | 9.31 | 55.4 | 51.8 | 59.9 | 55.7 | 66.9 | 53.3 | 46.9 | 55.7 | 94.4 | |||||
| Feather [8] | 20% | 9.31 | 54.3 | 54.8 | 59.8 | 54.8 | 65.1 | 53.6 | 45.6 | 55.9 | 94.7 | |||||
| FrameFusion [24] | 26% | - | 56.0 | 55.1 | 61.7 | 57.5 | 68.0 | 55.9 | 48.6 | 57.6 | 97.5 | |||||
| ShaRP | 23.5% | 9.31 | 57.1 | 56.1 | 62.4 | 58.0 | 68.5 | 56.1 | 48.6 | 58.4 | 98.9 | |||||
| FastV [6] | 15% | 7.66 | 54.9 | 54.8 | 58.4 | 54.7 | 65.1 | 53.4 | 45.7 | 55.7 | 94.4 | |||||
| PDrop [7] | 11.5% | 7.66 | 54.5 | 50.0 | 59.1 | 54.0 | 64.3 | 51.6 | 46.1 | 54.4 | 92.2 | |||||
| Feather [8] | 15% | 7.66 | 54.5 | 53.0 | 59.1 | 54.7 | 65.1 | 53.3 | 45.6 | 55.3 | 93.7 | |||||
| FrameFusion [24] | 21% | - | 57.0 | 54.3 | 60.6 | 56.6 | 67.0 | 55.2 | 47.6 | 57.1 | 96.8 | |||||
| ShaRP | 18.8% | 7.66 | 56.3 | 55.7 | 61.6 | 57.4 | 68.1 | 56.7 | 47.5 | 57.8 | 97.9 | |||||
| FastV [6] | 10% | 6.03 | 53.5 | 52.4 | 56.2 | 52.6 | 61.1 | 51.6 | 45.2 | 53.7 | 90.9 | |||||
| PDrop [7] | 1.5% | 6.03 | 48.2 | 44.7 | 52.5 | 45.0 | 49.6 | 45.8 | 39.6 | 47.6 | 80.6 | |||||
| Feather [8] | 10% | 6.03 | 53.6 | 52.7 | 57.6 | 53.6 | 62.4 | 52.7 | 45.6 | 54.4 | 92.1 | |||||
| FrameFusion [24] | 16% | - | 55.1 | 53.0 | 58.3 | 55.5 | 65.8 | 54.1 | 46.7 | 55.0 | 94.0 | |||||
| ShaRP | 14.0% | 6.03 | 55.9 | 55.8 | 61.4 | 56.4 | 65.9 | 56.3 | 46.9 | 57.4 | 97.2 | |||||
| Method |
MVBench ↑ |
VideoMME ↑ | Average ↑ | ||||||||||
| Overall | Short | Medium | Long | ||||||||||||||
| Duration | 16s | 160min | 160min | 13min | 330min | 3060min | Score | % | |||||||||
| Qwen2-VL [4] | 2880 | 100% | 16.7 | 64.8 | 55.1 | 56.7 | 67.4 | 54.2 | 48.3 | 58.9 | 100 | ||||||
| FastV [6] | 2880/576 | 20% | 4.1 | 57.7 | 48.0 | 52.6 | 61.9 | 49.2 | 46.7 | 52.8 | 89.6 | ||||||
| PDrop [7] | 2880/662 | 23.1% | 4.1 | 54.4 | 48.2 | 52.0 | 61.4 | 48.9 | 45.8 | 51.5 | 87.6 | ||||||
| Feather [8] | 2880/576 | 20% | 4.1 | 60.2 | 50.0 | 53.0 | 62.1 | 50.6 | 46.4 | 54.4 | 92.4 | ||||||
| FrameFusion [24] | 2880/740 | 25.7% | - | 60.4 | 48.5 | 53.4 | 62.4 | 50.7 | 47.2 | 54.1 | 91.9 | ||||||
| ShaRP | 2880/679 | 23.6% | 4.1 | 64.5 | 52.4 | 53.0 | 62.4 | 50.2 | 46.5 | 56.6 | 96.2 | ||||||
| FastV [6] | 2880/288 | 10% | 2.6 | 52.2 | 45.6 | 50.9 | 57.4 | 48.4 | 46.7 | 49.6 | 84.2 | ||||||
| PDrop [7] | 2880/78 | 3.4% | 2.6 | 46.1 | 45.3 | 46.6 | 52.8 | 44.3 | 42.6 | 46.0 | 78.1 | ||||||
| Feather [8] | 2880/288 | 10% | 2.6 | 56.8 | 48.6 | 51.1 | 60.2 | 47.9 | 45.1 | 52.2 | 88.6 | ||||||
| FrameFusion [24] | 2880/472 | 16.4% | - | 59.5 | 47.4 | 52.0 | 61.3 | 49.0 | 45.8 | 53.0 | 90.0 | ||||||
| ShaRP | 2880/400 | 13.9% | 2.6 | 63.9 | 53.3 | 52.2 | 61.5 | 49.1 | 46.1 | 56.5 | 95.9 | ||||||

### v2

# ShaRP: SHAllow-LayeR Pruning for Efficient Video Large Language Models

@@ Abstract
Video Large Language Models (VLLMs) incur substantial prefilling cost due to the large number of visual tokens. While attention-based token pruning offers a promising acceleration strategy, applying it at shallow decoder layers often causes severe performance degradation under high compression ratios, limiting its practical benefits. In this work, we uncover an overlooked failure mode in shallow-layer attention pruning: attention scores in early decoder layers can become unreliable indicators of token utility, resulting in unstable token selection under aggressive compression. We show that this effect arises from the joint influence of insufficient token interaction, content-agnostic positional bias, and redundancy among high-attention tokens, which together distort attention-based importance estimation before informative representations fully emerge. Motivated by this insight, we propose ShaRP, a unified pruning framework that restores reliable attention-based token selection by jointly improving local information aggregation, calibrating positional bias, and reducing redundancy. Extensive evaluations show that ShaRP preserves about 97.2% of the original performance while reducing TFLOPs by 86% and achieving a 5.1 speedup in the prefilling stage, providing a scalable solution for efficient VLLM inference, with no training.
@@ 1 Introduction
Vision Large Language Models (VLLMs) [1, 2, 3, 4, 5, 6] have brought deep video understanding task into a new era. They use a vision encoder to convert video streams into many tokens, which interact during pre-filling. Such a process thus incurs quadratic computational complexity relative to the token count. To mitigate this, numerous studies [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26] have focused on reducing visual token counts to compress information and accelerate VLLM inference.
Among various compression strategies, attention-based ones [7, 8, 9, 10, 27, 16] offer a compelling approach. Unlike visual cue-guided pruning that operates purely on visual similarity [12, 28, 29, 30, 20, 15, 26], attention-based methods leverage the inherent cross-modal reasoning ability of the model to estimate token importance, inherently aligning the pruning process with semantic relevance. It not only enhances interpretability [31, 32, 33] but also enables adaptive compression that dynamically focuses on content most relevant to the textual query. Representative studies such as FastV [7] show that visual information becomes less critical in deeper decoder layers, motivating aggressive pruning early. However, this strategy struggles with high compression on long videos, as performance drops sharply, leading subsequent works [8, 10, 22] to prune more conservatively in deeper layers, losing much of the computational savings. Viewed differently, we advocate the potential of attention in shallow layers as an effective criterion for token selection. To explore this, we conduct a pilot study comparing attention scores across shallow and deep layers. As shown in Fig. 2, visual analysis reveals that visual tokens near the sequence end receive abnormally high attention in shallow layers. When properly tackled, shallow layer attention exhibits distributions similar to deep ones, enabling early yet reliable pruning.
@@
To this end, we propose ShaRP, a training-free shallow-layer pruning framework built on a single insight: in early decoder layers, raw cross-attention is a miscalibrated importance signal. ShaRP restores a selective importance ranking that favors informative and non-redundant tokens via a lightweight three-step calibration loop. First, to enhance information interaction, inspired by [40], which observed that locally focused heads in Transformers form near-diagonal block attention, segmenting complex reasoning into phrasal units. We adapt this to the visual domain by partitioning videos into content-aware segments, intensifying local interactions and alleviating attention collapse [36, 37]. Second, our method calibrates positional bias in attention scores through a rapid global bias estimation. Finally, to balance between the importance and redundancy, high-attention register tokens are refined through a three-step pipeline composed of pre-filtering, deduplication, and post-filling, which eliminates redundancies while enriching the representation with broader contextual information. Extensive experiments demonstrate that our approach achieves high compression in shallow LLM layers while maintaining state-of-the-art performance, establishing a new paradigm for attention-based token pruning.
@@ 2.1 Visual Cue-Guided Token Pruning
Since their emergence, VLLMs [41, 42, 43, 2, 44, 45, 4] have demonstrated significant potential across diverse applications, including VQA [46, 47, 48], video captioning [49, 50], contextual reasoning [51, 52, 53] etc.. Earlier approaches employ a limited number of learnable queries to aggregate visual information. For instance, VideoLLaMA [54] adopts a Q-Former [45] to distill dense visual features into compact query sets. While GlimpsePrune [55] introduces a trainable Visual Importance Predictor (VIP) within the VLLM to score and prune visual tokens at different layers. The recent method EPIC [56] adopts a Progressive Consistency Distillation strategy to re-finetune the model, enabling it to learn rich visual information even from incomplete visual inputs. Token pruning can be achieved in a training-free manner. Among these, LLaVA-PruMerge [11], VisionZip [12], and VisPruner [14] leverage the inherent sensitivity of vision encoders to visual tokens to identify and retain visually dominant tokens. DivPrune [28] formulates pruning as a Max-Min Diversity Problem (MMDP), ensuring that the set of retained visual tokens possesses maximum diversity. Nevertheless, the temporal redundancy inherent in video data calls for temporal-aware pruning mechanisms. DyCoke [15] merges tokens within fixed temporal windows. Other methods [29, 20, 21, 22] remap the numerous visual tokens back to the frame level, enabling them to identify and merge highly similar visual content across frames while retaining unique content. Among these, STTM [20] performs multi-granular spatio-temporal token merging. FastVID [21] performs segmentation based on video content, combining temporal segmentation with spatio-temporal merging. HoliTom [22] utilizes dynamic programming to merge redundant information across adjacent frames.
@@ 4.2 Positional Bias Calibration
Positional Encodings (PE) equip VLLMs with token order information, whereas also introduce inherent positional bias [59] that favors tokens at specific locations. A typical example is the long-term decay issue [36] in RoPE (Eq. 2), where attention scores decrease as the relative token distance increases. While in the context of video token pruning, existing methods [7, 8, 10] usually identify more important visual tokens by referring attention scores from the last text token . They typically assume that visual tokens with higher scores are more relevant to the query.
However, we argue that positional bias significantly distorts the distribution of attention scores, which may lead to misinterpretation of this mechanism. As shown in Fig. 4 (left-top), our analysis on the VideoMME [58] benchmark at 20% retention reveals that most top-ranked tokens (orange regions) cluster near the final frames, reflecting a clear positional bias rather than true semantic relevance. Fig. 4 (right-middle) further illustrates a case where the question targets a mid-video event, yet attention remains skewed toward later frames, leaving semantically important regions under-attended (blue regions).
@@ 4.3 Pruning and Register Token Deduplication
| Method | Retention Ratio ↓ | TFLOPs ↓ | MVBench ↑ | LongVideo Bench ↑ | MLVU ↑ | VideoMME ↑ | Average ↑ | ||||
| Overall | Short | Medium | Long | ||||||||
| Duration | 16s | 160min | 3120min | 160min | 13min | 330min | 3060min | Score | % | ||
[…]
| FastV [7] | 25% | 11 | 55.3 | 55.1 | 60.3 | 56.5 | 68.2 | 55 | 46.3 | 56.8 | 96.2 |
| PDrop [8] | 28.5% | 11 | 56.7 | 54.2 | 61.1 | 56.4 | 68.0 | 53.6 | 47.7 | 57.1 | 96.7 |
| Feather [10] | 25% | 11 | 54.8 | 54.7 | 60.8 | 55.0 | 65.3 | 53.6 | 46.2 | 56.3 | 95.4 |
| Dycoke [15] | 25% | 8.7 | 53.1 | 49.5 | 55.8 | 51.0 | 61.1 | 48.6 | 43.2 | 52.4 | 88.7 |
| STTM [20] | 21.2% | 10.5 | 50.8 | 51.1 | 58.7 | 54.9 | 65.2 | 53.5 | 46.1 | 53.9 | 91.3 |
| FrameFusion [30] | 31% | - | 56.7 | 55.8 | 62.2 | 57.4 | 68.1 | 55.4 | 48.6 | 57.8 | 97.9 |
| ShaRP | 28.2% | 11 | 57.1 | 56.9 | 62.6 | 57.7 | 68.4 | 55.7 | 48.9 | 58.6 | 99.3 |
| FastV [7] | 20% | 9.31 | 55.8 | 55.3 | 59.5 | 55.8 | 66.4 | 54.2 | 46.8 | 56.6 | 95.9 |
| PDrop [8] | 19.6% | 9.31 | 55.4 | 51.8 | 59.9 | 55.7 | 66.9 | 53.3 | 46.9 | 55.7 | 94.4 |
| Feather [10] | 20% | 9.31 | 54.3 | 54.8 | 59.8 | 54.8 | 65.1 | 53.6 | 45.6 | 55.9 | 94.7 |
| FrameFusion [30] | 26% | - | 56.0 | 55.1 | 61.7 | 57.5 | 68.0 | 55.9 | 48.6 | 57.6 | 97.5 |
| ShaRP | 23.5% | 9.31 | 57.1 | 55.8 | 61.9 | 57.6 | 68.1 | 56.1 | 48.6 | 58.1 | 98.5 |
| FastV [7] | 15% | 7.66 | 54.9 | 54.8 | 58.4 | 54.7 | 65.1 | 53.4 | 45.7 | 55.7 | 94.4 |
| PDrop [8] | 11.5% | 7.66 | 54.5 | 50.0 | 59.1 | 54.0 | 64.3 | 51.6 | 46.1 | 54.4 | 92.2 |
| Feather [10] | 15% | 7.66 | 54.5 | 53.0 | 59.1 | 54.7 | 65.1 | 53.3 | 45.6 | 55.3 | 93.7 |
| FrameFusion [30] | 21% | - | 57.0 | 54.3 | 60.6 | 56.6 | 67.0 | 55.2 | 47.6 | 57.1 | 96.8 |
| ShaRP | 18.8% | 7.66 | 56.5 | 55.9 | 62.5 | 56.3 | 67.8 | 53.5 | 47.7 | 57.8 | 98.0 |
| FastV [7] | 10% | 6.03 | 53.5 | 52.4 | 56.2 | 52.6 | 61.1 | 51.6 | 45.2 | 53.7 | 90.9 |
| PDrop [8] | 1.5% | 6.03 | 48.2 | 44.7 | 52.5 | 45.0 | 49.6 | 45.8 | 39.6 | 47.6 | 80.6 |
| Feather [10] | 10% | 6.03 | 53.6 | 52.7 | 57.6 | 53.6 | 62.4 | 52.7 | 45.6 | 54.4 | 92.1 |
| FrameFusion [30] | 16% | - | 55.1 | 53.0 | 58.3 | 55.5 | 65.8 | 54.1 | 46.7 | 55.0 | 94.0 |
| ShaRP | 14.1% | 6.03 | 56.4 | 55.3 | 62.0 | 55.8 | 65.9 | 54.6 | 46.9 | 57.4 | 97.2 |
| Method | Retention Ratio ↓ | TFLOPs ↓ | MVBench ↑ | LongVideo Bench ↑ | VideoMME ↑ | Average ↑ | ||||
| Overall | Short | Medium | Long | |||||||
| Duration | 16s | 160min | 160min | 13min | 330min | 3060min | Score | % | ||
[…]
| FastV [7] | 20.0% | 16.3 | 54.5 | 54.6 | 58.7 | 68.8 | 58.5 | 48.8 | 55.9 | 91.9 |
| PDrop [8] | 19.5% | 16.3 | 55.1 | 55.0 | 60.3 | 70.6 | 60.4 | 49.9 | 56.8 | 93.4 |
| Fearher [10] | 20.0% | 16.3 | 54.0 | 55.0 | 59.7 | 71.3 | 58.2 | 49.5 | 56.2 | 92.4 |
| Dycoke [15] | 27.3% | 16.3 | 51.3 | 53.5 | 56.6 | 66.3 | 54.1 | 49.4 | 53.8 | 88.5 |
| STTM [20] | 17.1% | 17.0 | 49.2 | 54.5 | 59.8 | 68.7 | 59.8 | 50.9 | 54.5 | 89.6 |
| Framefusion [30] | 25.5% | - | 56.5 | 57.6 | 61.3 | 72.8 | 59.3 | 51.9 | 58.5 | 96.2 |
| ShaRP | 23.5% | 16.3 | 58.2 | 58.1 | 62.7 | 75.7 | 60.9 | 52.5 | 59.7 | 98.2 |
[…]
| FastV [7] | 10% | 2.2 | 52.2 | 45.6 | 50.9 | 57.4 | 48.4 | 46.7 | 49.6 | 84.2 |
| PDrop [8] | 3.4% | 2.2 | 46.1 | 45.3 | 46.6 | 52.8 | 44.3 | 42.6 | 46.0 | 78.1 |
| Feather [10] | 10% | 2.2 | 56.8 | 48.6 | 51.1 | 60.2 | 47.9 | 45.1 | 52.2 | 88.6 |
| FrameFusion [30] | 16.4% | - | 59.5 | 47.4 | 52.0 | 61.3 | 49.0 | 45.8 | 53.0 | 90.0 |
| ShaRP | 13.9% | 2.2 | 63.9 | 53.3 | 52.2 | 61.5 | 49.1 | 46.1 | 56.5 | 95.9 |
[…]
| FastV [7] | 10.0% | 3.4 | 60.2 | 47.7 | 54.6 | 61.5 | 55.0 | 47.3 | 54.2 | 86.6 |
| PDrop [8] | 3.4% | 3.4 | 56.0 | 47.4 | 52.4 | 54.4 | 52.6 | 49.3 | 51.9 | 83.0 |
| Fearher [10] | 10.0% | 3.4 | 61.6 | 50.5 | 55.0 | 53.3 | 53.9 | 47.8 | 55.7 | 89.0 |
| Framefusion [30] | 16.4% | - | 63.3 | 50.4 | 56.4 | 56.4 | 66.0 | 55.1 | 56.7 | 90.6 |
| ShaRP | 14.1% | 3.4 | 63.8 | 53.9 | 56.8 | 66.3 | 55.6 | 48.3 | 58.2 | 93.0 |